OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = 625 and a3 = 25?

OpenStudy (anonymous):

\[r =\frac{1}{5}\]

OpenStudy (anonymous):

thank you satellite :)

OpenStudy (anonymous):

so it will be \[625(\frac{1}{5})^5\]

OpenStudy (anonymous):

or we could write \[a_3=25\] \[a_4=5\] \[a_5=1\] \[a_6=\frac{1}{5}\]

OpenStudy (anonymous):

\[\[a_3=25\] \[a_4=5\] \[a_5=1\] \[a_6=\frac{1}{5}\] \]